why is it impossible to draw an isosceles triangle with one angle that measures 52 degrees and a second angle that measures 74 degrees

Answer by stanbon(75887)

Why is it impossible to draw an isosceles triangle with
one angle that measures 52 degrees and a second angle
that measures 74 degrees.
-----------------
Isosceles implies there are 2 equal angles.
---
If the equal angles are each 52 degrees,
2*52+74 = 104 + 74 = 178 degrees
---
So, it is not impossible.
-------------
Cheers,
Stan H.
-------------